# Unbiased Sliced Wasserstein Kernels for High-Quality Audio Captioning

### v1

@@ Abstract
Teacher-forcing training for audio captioning usually leads to exposure bias due to training and inference mismatch. Prior works propose the contrastive method to deal with caption degeneration. However, the contrastive method ignores the temporal information when measuring similarity across acoustic and linguistic modalities, leading to inferior performance. In this work, we develop the temporal-similarity score by introducing the unbiased sliced Wasserstein RBF (USW-RBF) kernel equipped with rotary positional embedding to account for temporal information across modalities. In contrast to the conventional sliced Wasserstein RBF kernel, we can form an unbiased estimation of USW-RBF kernel via Monte Carlo estimation. Therefore, it is well-suited to stochastic gradient optimization algorithms, and its approximation error decreases at a parametric rate of $\mathcal{O}(L^{-1/2})$ with $L$ Monte Carlo samples. Additionally, we introduce an audio captioning framework based on the unbiased sliced Wasserstein kernel, incorporating stochastic decoding methods to mitigate caption degeneration during the generation process. We conduct extensive quantitative and qualitative experiments on two datasets, AudioCaps and Clotho, to illustrate the capability of generating high-quality audio captions. Experimental results show that our framework is able to increase caption length, lexical diversity, and text-to-audio self-retrieval accuracy.
Machine Learning, ICML
[…]
tcb@breakable
@@ 1 Introduction
Dynamic Time Warping (DTW(Sakoe & Chiba, 1978) and  Soft Dynamic Time Warping (soft-DTW(Cuturi & Blondel, 2017) are two widely adopted distances used to measure the discrepancy between two time series. They are capable of considering temporal information, however, the monotonic alignment imposed by DTW is too strict and might adversely affect the measurement of the discrepancy between audio and caption when local temporal distortion exists. (Su & Hua, 2017) proposed an order-preserving Wasserstein distance to deal with the shortcoming of DTW. Although the order-preserving Wasserstein distance can measure the discrepancy between two sequential data when temporal distortion exists, it is ineffective to measure the discrepancy between high-dimensional sequences due to the dimensionality curse of the Wasserstein distance.
To address all aforementioned issues, we propose the Audio Captioing with Unbiased sliced Wasserstein kernel (ACUS)framework to alleviate the caption degeneration for the audio captioning task and better measure cross-modal similarity. We develop the unbiased sliced Wasserstein RBF kernel (USW-RBF)for precisely measuring the similarity score between acoustic and linguistic modalities. The USW-RBF leverages the radial basis function (RBF) kernel, in which the sliced Wasserstein distance equipped with the rotary positional embedding is used as the distance. The proposed kernel is unbiased. Hence, it is highly compatible with stochastic gradient optimization algorithms, and its approximation error decreases at a parametric rate of $\mathcal{O}(L^{-1/2})$. We also derive the proposed kernel and show that it is capable of measuring the similarity in terms of features and temporal information. Furthermore,  (Arora et al., 2022a) provides an analysis of exposure bias through the lens of imitation learning and empirically shows that stochastic decoding methods are able to alleviate exposure bias for language models. According to this observation, we leverage the ACUS framework with stochastic decoding methods at the inference stage to rerank generated captions to choose the most suitable candidate caption. To sum up, our contributions can be summarized as follows:
@@
    We propose the USW-RBF kernel to precisely measure the similarity between acoustic and linguistic modalities for encoder-decoder audio captioning models. Our kernel is able to deal with the dimensionality curse and temporal distortion by leveraging the sliced Wasserstein distance equipped with rotary positional embedding.
@@ 2.1 Encoder-Decoder Audio Captioning
An encoder-decoder audio captioning model, denoted as $\mathcal{M}=(f_{\theta},g_{\phi})$, is capable of generating captions $\mathbf{y}=\{y_{t}\}_{t=0}^{N}$ conditioning on a given audio $\mathbf{x}$. Here, $f_{\theta}$ ($\theta\in\Theta$) and $g_{\phi}$ ($\phi\in\Phi$) are the encoder and decoder parameterized by $\theta$ and $\phi$ respectively. The encoder is designed to extract acoustic features from audio, while the decoder is able to decode extracted acoustic features to natural language. The audio captioning model is trained to maximize the likelihood of ground-truth captions when predicting the current word in the sequence given the prior words $y_{<t}$ and the hidden representation of audio $z_{x}=f_{\theta}(x)$. The training objective for the audio captioning model is defined as follows:
$\mathcal{L}_{MLE}=-\sum_{t=1}^{N}\log p_{g_{\phi}}(y_{t}|z_{x},y_{<t}).$ (1)
@@
$\mathbf{\hat{y}}=\operatorname*{arg\,max}_{\mathbf{y_{i}}\in\mathcal{B}}p_{g_{\phi}}(\mathbf{y_{i}}|z_{x}).$ (2)
There is a critical issue with likelihood training, which is exposure bias. The audio captioning model predicts the next word based on previous ground-truth words $y_{<t}\in y$ at the training stage, but it adopts the predicted tokens $\hat{y}_{<t}$ by itself to generate the next token $\hat{y_{t}}$ at inference stage. Due to exposure bias, there is a significant gap in terms of performance of pretrained audio captioning models on training and test data. Furthermore, the beam search decoding even makes the exposure bias more critical due to error accumulation.
@@ 2.2 Contrastive Learning for Audio Captioning
$\mathcal{L}_{NCE}=-\log\frac{\exp(\cos(z_{x},z_{y})/\tau)}{\sum_{y^{\prime}\in Y}\exp(\cos(z_{x},z_{y^{\prime}})/\tau)},$ (3)
where $z_{x},z_{y},z_{y^{\prime}}\in\mathbb{R}^{d}$ denote the hidden representation of audio input $x$, ground-truth caption $y$, and caption $y^{\prime}\in Y$ from the minibatch, respectively. The temperature $\tau>0$ is utilized to control the strength of penalties on negative examples. The likelihood objective is jointly optimized with the contrastive loss at the training phase
@@
$\mathbf{\hat{y}}=\operatorname*{arg\,max}_{\mathbf{y_{i}}\in\mathcal{B}}\{p_{g_{\theta}}(\mathbf{y_{i}}|z_{x})+\cos(z_{x},z_{y_{i}})\}.$ (5)
Although contrastive regularization is effective in mitigating exposure bias for audio captioning, the similarity between audio and ground-truth caption hidden representation is computed based on cosine similarity between the average pooling of audio and caption hidden representation. The average pooling operation discards the temporal information in audio and caption representation, therefore, leveraging contrastive regularization for inference can lead to inferior performance.
@@ 3 Methodology
We first develop the unbiased sliced Wasserstein RBF kernel (USW-RBF)to deal with the dimensionality curse and strict monotonic alignment for measuring similarity across multimodalities. The USW-RBF is equipped with the rotary positional embedding to consider temporal information when measuring similarity across linguistic and acoustic modalities. Then, we propose the Audio Captioing with Unbiased sliced Wasserstein kernel (ACUS)framework to mitigate text degeneration for audio captioning. We leverage stochastic decoding methods with the USW-RBF as similarity score across modality to alleviate exposure bias at the inference stage. Our training and inference procedures are illustrated in Figure 1.
@@ 3.1 Unbiased Sliced Wasserstein Kernel
Wasserstein distance. Given $p\geq 1$, Wasserstein distance (Peyré et al., 2019) between $\mu$ and $\nu$ be two distributions belongs to $\mathcal{P}_{p}(\mathbb{R}^{d})$ is defined as:
[…]
where $\Pi(\mu,\nu)$ is the set of all distributions that has the first marginal is $\mu$ and the second marginal is $\nu$ i.e., transportation plans or couplings.
Sliced Wasserstein distance. Given $p\geq 1$, the sliced Wasserstein (SW) distance (Bonneel et al., 2015; Nguyen et al., 2021; Nguyen & Ho, 2024) between two probability distributions $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ is defined as:
@@
[…]
where $F_{\psi\sharp\mu}$ and $F_{\psi\sharp\nu}$ are the cumulative distribution function (CDF) of $\psi\sharp\mu$ and $\psi\sharp\nu$ respectively. When $\mu$ and $\nu$ are empirical distributions over sets $Z_{x}=\{z_{x}^{1},\ldots,z_{x}^{N}\}$ and $Z_{y}=\{z_{y}^{1},\ldots,z_{y}^{M}\}$ i.e., $\mu=\frac{1}{N}\sum_{i=1}^{N}\delta_{z_{x}^{i}}$ and $\nu=\frac{1}{M}\sum_{j=0}^{M}\delta_{z_{y}^{j}}$ respectively, $\psi\sharp\mu$ and $\psi\sharp\nu$ are empirical distributions over sets $\psi^{\top}Z_{x}=\{\psi^{\top}z_{x}^{1},\ldots,\psi^{\top}z_{x}^{N}\}$ and $\psi^{\top}Z_{y}=\{\psi^{\top}z_{y}^{1},\ldots,\psi^{\top}z_{y}^{M}\}$ in turn (by abusing the notation of matrix multiplication). As a result, the quantile functions can be approximated efficiently.
@@
where $L$ is referred to as the number of projections. When two empirical distributions have the same number of supports i.e., $\mu=\frac{1}{N}\sum_{i=1}^{N}\delta_{z_{x}^{i}}$ and $\nu=\frac{1}{N}\sum_{j=0}^{N}\delta_{z_{y}^{j}}$, we have:
$\widehat{SW}^{p}_{p}(\mu,\nu;L)=\frac{1}{L}\frac{1}{N}\sum_{l=1}^{L}\sum_{i=1}^{N}\|\psi^{\top}z_{x}^{\sigma_{1,l}(i)}-\psi^{\top}z_{y}^{\sigma_{2,l}(i)}\|_{p}^{p},$
where $\sigma_{1,l}:[[N]]\to[[N]]$ and $\sigma_{2,l}:[[N]]\to[[N]]$ are two sorted permutation mapping of $\psi^{\top}Z_{x}$ and $\psi^{\top}Z_{y}$ in turn. By abusing of notation, we will use the notation $\widehat{SW}^{p}_{p}(Z_{x},Z_{y};L)$ later when $\mu$ and $\nu$ are empirical distributions over $Z_{x}$ and $Z_{y}$.
Sliced Wasserstein RBF kernels. Given the definition of SW in Equation (6), we can define the sliced Wasserstein RBF (SW-RBF) kernel (Carriere et al., 2017; Kolouri et al., 2016) as:
@@
Definition 3.1.
Given two probability distributions $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$, $\kappa\in\mathbb{R}_{+}$, $p\geq 1$, the unbiased sliced Wasserstein RBF kernel (USW-RBF) is defined as:
@@
Proof of Proposition 3.2 is given in Appendix A.1.1. Since the USW-RBF kernel is positive definite, it is equivalent to a reproducing kernel Hilbert space and celebrates the representer theorem.
Proposition 3.3.
@@
It is worth noting that Quasi-Monte Carlo methods (Nguyen et al., 2024) and control variates techniques (Nguyen & Ho, 2023; Leluc et al., 2024) can also be applied to achieve more accurate approximation. However, we use the basic Monte Carlo to make theoretical investigation easier.
Proposition 3.4.
@@
(i) $\widehat{\mathcal{UK}}_{\gamma}(\mu,\nu;p,L)$ is an unbiased estimate of $\mathcal{UK}_{\gamma}(\mu,\nu)$ i.e., $\mathbb{E}[\widehat{\mathcal{UK}}_{\gamma}(\mu,\nu;p,L)]=\mathcal{UK}_{\gamma}(\mu,\nu;p)$,
@@
The proof of Proposition 3.4 is given in Appendix A.1.3. The unbiasedness (i) is crucial for the convergence of stochastic gradient algorithms, which optimizes the kernel as a loss. The bound in (ii) suggests that the approximation error decreases at a parametric rate of $\mathcal{O}(L^{-1/2})$.

### v2

@@ Abstract
Audio captioning systems face a fundamental challenge: teacher-forcing training creates exposure bias that leads to caption degeneration during inference. While contrastive methods have been proposed as solutions, they typically fail to capture the crucial temporal relationships between acoustic and linguistic modalities. We address this limitation by introducing the unbiased sliced Wasserstein RBF (USW-RBF) kernel with rotary positional embedding, specifically designed to preserve temporal information across modalities. Our approach offers a practical advantage: the kernel enables efficient stochastic gradient optimization, making it computationally feasible for real-world applications. Building on this foundation, we develop a complete audio captioning framework that integrates stochastic decoding to further mitigate caption degeneration. Extensive experiments on AudioCaps and Clotho datasets demonstrate that our method significantly improves caption quality, lexical diversity, and text-to-audio retrieval accuracy. Furthermore, we demonstrate the generalizability of our USW-RBF kernel by applying it to audio reasoning tasks, where it enhances the reasoning capabilities of large audio language models on the CompA-R in terms of correctness and quality. Our kernel also improves the reasoning accuracy of the MMAU-test-mini benchmarks by $4\%$. These results establish our approach as a powerful and generalizable solution for cross-modal alignment challenges in audio-language tasks.
@@ 1 Introduction
Dynamic Time Warping (DTW) (Sakoe and Chiba, 1978) and  Soft Dynamic Time Warping (soft-DTW) (Cuturi and Blondel, 2017) are two widely adopted distances used to measure the discrepancy between two time series. They are capable of considering temporal information, however, the monotonic alignment imposed by DTW is too strict and might adversely affect the measurement of the discrepancy between audio and caption when local temporal distortion exists. (Su and Hua, 2017) proposed an order-preserving Wasserstein distance to deal with the shortcoming of DTW. Although the order-preserving Wasserstein distance can measure the discrepancy between two sequential data when temporal distortion exists, it is ineffective to measure the discrepancy between high-dimensional sequences due to the dimensionality curse of the Wasserstein distance.
To address all aforementioned issues, we propose the Audio Captioning with Unbiased sliced Wasserstein kernel (ACUS) framework to alleviate the caption degeneration for the audio captioning task and better measure cross-modal similarity. We develop the unbiased sliced Wasserstein RBF kernel (USW-RBF) for precisely measuring the similarity score between acoustic and linguistic modalities. The USW-RBF leverages the radial basis function (RBF) kernel, in which the sliced Wasserstein distance equipped with the rotary positional embedding is used as the distance. The proposed kernel is unbiased. Hence, it is highly compatible with stochastic gradient optimization algorithms (Dai et al., 2014), and its approximation error decreases at a parametric rate of $\mathcal{O}(L^{-1/2})$. We also derive the proposed kernel and show that it is capable of measuring the similarity in terms of features and temporal information. Furthermore,  (Arora et al., 2022a) provides an analysis of exposure bias through the lens of imitation learning and empirically shows that stochastic decoding methods are able to alleviate exposure bias for language models. According to this observation, we leverage the ACUS framework with stochastic decoding methods at the inference stage to rerank generated captions to choose the most suitable candidate caption. Our contributions can be summarized as follows:
@@
    We propose the USW-RBF kernel to precisely measure the similarity between acoustic and linguistic modalities for encoder-decoder audio captioning models. Our kernel is able to deal with temporal distortion by leveraging the sliced Wasserstein distance equipped with rotary positional embedding. The experimental results from audio captioning and reasoning tasks demonstrate the ability of our kernel to measure cross-modal alignment between acoustic and linguistic modalities.
@@ 2.1 Encoder-Decoder Audio Captioning
An encoder-decoder audio captioning model, denoted as $\mathcal{M}=(f_{\theta},g_{\phi})$, is capable of generating captions $\mathbf{y}=\{y_{t}\}_{t=0}^{N}$ conditioning on a given audio $\mathbf{x}$. Here, $f_{\theta}$ and $g_{\phi}$ are the encoder and decoder parameterized by $\theta$ and $\phi$ respectively. The encoder is designed to extract acoustic features from audio, while the decoder is able to decode extracted acoustic features to natural language. The audio captioning model is trained to maximize the likelihood of ground-truth captions when predicting the current word in the sequence given the prior words $y_{<t}$ and the hidden representation of audio $z_{\mathbf{x}}=f_{\theta}(\mathbf{x})$. The training objective for the audio captioning model is defined as follows:
$\mathcal{L}_{MLE}=-\sum_{t=1}^{N}\log p_{g_{\phi}}(y_{t}|z_{\mathbf{x}},y_{<t}).$ (1)
@@
$\mathbf{\hat{y}}=\operatorname*{arg\,max}_{\mathbf{y}\in\mathcal{B}}p_{g_{\phi}}(\mathbf{y}|z_{\mathbf{x}}).$ (2)
Limitation of likelihood training. There is a critical issue with likelihood training, which is exposure bias. The audio captioning model predicts the next word based on previous ground-truth words $y_{<t}\in\mathbf{y}$ at the training stage, but it adopts the predicted tokens $\hat{y}_{<t}$ by itself to generate the next token $\hat{y_{t}}$ at inference stage. Due to exposure bias, there is a significant gap in terms of performance of pretrained audio captioning models on training and test data. Furthermore, the beam search decoding even makes the exposure bias more critical due to error accumulation.
@@ 2.2 Contrastive Learning for Audio Captioning
$\mathcal{L}_{NCE}=-\log\frac{\exp(\cos(z_{\mathbf{x}},z_{\mathbf{y}})/\tau)}{\sum_{\mathbf{y}^{\prime}\in Y}\exp(\cos(z_{\mathbf{x}},z_{\mathbf{y}^{\prime}})/\tau)},$ (3)
where $z_{\mathbf{x}},z_{\mathbf{y}},z_{\mathbf{y}^{\prime}}\in\mathbb{R}^{d}$ denote the hidden representation of audio input $\mathbf{x}$, the ground-truth caption $\mathbf{y}$, and the caption $\mathbf{y}^{\prime}$ from the minibatch of captions $Y$, respectively. The temperature $\tau>0$ is utilized to control the strength of penalties on negative examples. The likelihood objective is jointly optimized with the contrastive loss at the training phase
@@
$\mathbf{\hat{y}}=\operatorname*{arg\,max}_{\mathbf{y}\in\mathcal{B}}\{p_{g_{\phi}}(\mathbf{y}|z_{\mathbf{x}})+\cos(z_{\mathbf{x}},z_{\mathbf{y}})\}.$ (5)
Limitation of contrastive learning. Although contrastive regularization is effective in mitigating exposure bias for audio captioning, the cross-modal alignment between acoustic and linguistic modalities is computed based on the cosine similarity between either the average pooling or weighted aggregation of audio and caption hidden representations. These aggregation methods discard the temporal information in audio and caption representations, therefore, leveraging contrastive regularization for inference can lead to inferior performance.
@@ 3 Methodology
We first develop USW-RBF to deal with temporal distortion when measuring similarity across multimodalities. The USW-RBF is equipped with the rotary positional embedding to consider temporal information when measuring similarity across linguistic and acoustic modalities. Then, we propose the ACUS framework to mitigate text degeneration for audio captioning. We leverage stochastic decoding methods with the USW-RBF as a similarity score across modality to alleviate exposure bias at the inference stage. Our training and inference procedure are illustrated in Figure 1.
@@ 3.1 Unbiased Sliced Wasserstein Kernel
Wasserstein distance. Given $p\geq 1$, a Wasserstein distance (Peyré et al., 2019) between two distributions, $\mu$ and $\nu$ , in $\mathcal{P}_{p}(\mathbb{R}^{d})$ is defined as:
$\text{W}_{p}^{p}(\mu,\nu)$ $=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{p}d\pi(x,y)$
where $\Pi(\mu,\nu)$ is the set of all distributions that has the first marginal is $\mu$ and the second marginal is $\nu$ , i.e., transportation plans or couplings.
Sliced Wasserstein distance. Given $p\geq 1$, the sliced Wasserstein (SW) distance Bonneel et al. (2015); Nguyen et al. (2021); Nguyen and Ho (2024) between two probability distributions $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ is defined as:
@@
$\text{W}_{p}^{p}(\psi\sharp\mu,\psi\sharp\nu)=\int_{0}^{1}|F_{\psi\sharp\mu}^{-1}(z)-F_{\psi\sharp\nu}^{-1}(z)|^{p}dz$
where $\sharp$ denotes the push-forward projection, while $F_{\psi\sharp\mu}$ and $F_{\psi\sharp\nu}$ are the cumulative distribution function (CDF) of $\psi\sharp\mu$ and $\psi\sharp\nu$ respectively. When $\mu$ and $\nu$ are empirical distributions over sets $Z_{\mathbf{x}}=\{z_{\mathbf{x}}^{1},\ldots,z_{\mathbf{x}}^{N}\}$ and $Z_{\mathbf{y}}=\{z_{\mathbf{y}}^{1},\ldots,z_{\mathbf{y}}^{M}\}$, i.e., $\mu=\frac{1}{N}\sum_{i=1}^{N}\delta_{z_{\mathbf{x}}^{i}}$ and $\nu=\frac{1}{M}\sum_{j=0}^{M}\delta_{z_{\mathbf{y}}^{j}}$ respectively, $\psi\sharp\mu$ and $\psi\sharp\nu$ are empirical distributions over sets $\psi^{\top}Z_{\mathbf{x}}=\{\psi^{\top}z_{\mathbf{x}}^{1},\ldots,\psi^{\top}z_{\mathbf{x}}^{N}\}$ and $\psi^{\top}Z_{\mathbf{y}}=\{\psi^{\top}z_{\mathbf{y}}^{1},\ldots,\psi^{\top}z_{\mathbf{y}}^{M}\}$ in turn (by abusing the notation of matrix multiplication). As a result, the quantile functions can be approximated efficiently.
@@
where $L$ is referred to as the number of projections. When two empirical distributions have the same number of supports, i.e., $\mu=\frac{1}{N}\sum_{i=1}^{N}\delta_{z_{\mathbf{x}}^{i}}$ and $\nu=\frac{1}{N}\sum_{j=0}^{N}\delta_{z_{\mathbf{y}}^{j}}$, we have:
$\widehat{SW}^{p}_{p}(\mu,\nu;L)=\frac{1}{L}\frac{1}{N}\sum_{l=1}^{L}\sum_{i=1}^{N}\|\psi^{\top}z_{\mathbf{x}}^{\sigma_{1,l}(i)}-\psi^{\top}z_{\mathbf{y}}^{\sigma_{2,l}(i)}\|_{p}^{p},$
where $\sigma_{1,l}:[[N]]\to[[N]]$ and $\sigma_{2,l}:[[N]]\to[[N]]$ are two sorted permutation mapping of $\psi^{\top}Z_{\mathbf{x}}$ and $\psi^{\top}Z_{\mathbf{y}}$ in turn. By abusing of notation, we use the notation $\widehat{SW}^{p}_{p}(Z_{\mathbf{x}},Z_{\mathbf{y}};L)$ later when $\mu$ and $\nu$ are empirical distributions over $Z_{\mathbf{x}}$ and $Z_{\mathbf{y}}$.
Sliced Wasserstein RBF kernels. Given the definition of SW in Equation (6), the definition of sliced Wasserstein RBF (SW-RBF) kernel (Carriere et al., 2017; Kolouri et al., 2016) is:
@@
Definition 1.
Given two probability distributions $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$, $\gamma\in\mathbb{R}_{+}$, $p\geq 1$, the unbiased sliced Wasserstein RBF kernel (USW-RBF) is defined as:
@@
Proof of Proposition 1 is given in Appendix A.1.1. Since the USW-RBF kernel is positive definite, it is equivalent to a reproducing kernel Hilbert space and celebrates the representer theorem.
Proposition 2.
@@
It is worth noting that Quasi-Monte Carlo methods (Nguyen et al., 2024) and control variates techniques (Nguyen and Ho, 2023; Leluc et al., 2024) can also be applied to achieve more accurate approximation. However, we use the basic Monte Carlo to make theoretical investigation easier.
Proposition 3.
@@
(i) $\widehat{\mathcal{UK}}_{\gamma}(\mu,\nu;p,L)$ is an unbiased estimate of $\mathcal{UK}_{\gamma}(\mu,\nu;p)$ , i.e., $\mathbb{E}[\widehat{\mathcal{UK}}_{\gamma}(\mu,\nu;p,L)]=\mathcal{UK}_{\gamma}(\mu,\nu;p)$,
@@
The proof of Proposition 3 is given in Appendix A.1.3. The unbiasedness (i) is crucial for the convergence of stochastic gradient algorithms (Bottou et al., 2018) which optimizes the kernel as a loss. The bound in (ii) suggests that the approximation error decreases at a parametric rate of $\mathcal{O}(L^{-1/2})$.